Grade 11Algebra

solve log5120 + (x-3) - 2 * log5 (1-5(x-3)) = -log5 (0.2- 5(x-4))

radha suresh
11 Years agoGrade 11

1 Answer

Nicho priyatham
11 Years ago
bring all log terms to 1 side and put (x-3) on other side now 
now add all log terms  using the formulas
  •   mlogab=logabm
  • log a +log b =log(ab)
  • log a-log b =log(a/b)
and REMOVE log on both sides to get the equation
and put 5x-3=a to get the equation
(1-a)2=24a(1-a)
on solving v get a= 1,1/25
so x=3,1
BUT wn we put x=3 in expression we get log50 wh is not define so 
x=1 is oniy solution
PLZ APPROVE IF USEFULL  incase of any doubts drop them in ans box